OpenStudy (anonymous):

Solve these systems of equations 7x+2y= -3 and -14x-4y=6 . I got y=3/2, x= -6/7 am I right ?

OpenStudy (aravindg):

hmm multiply first equation with 2

OpenStudy (aravindg):

14x+4y=-6

OpenStudy (aravindg):

14x+4y+6=0

OpenStudy (anonymous):

well yeah i did that multiplying it by -2 so that when i add it in the second question -14x and 14X cancel out (I'm using linear combinations here)

OpenStudy (aravindg):

-14x+4y-6=0

OpenStudy (anonymous):

*equation

OpenStudy (aravindg):

if thats the case both x and y cancels out!!

OpenStudy (aravindg):

isnt it?

OpenStudy (anonymous):

In the first equation you do 2y times -2 = -4y. Now in the 2nd equation you have -4y soo -4y+ -4y is -8y which means that the ys don't cancel out

OpenStudy (aravindg):

nt really :)

OpenStudy (aravindg):

infact both these lines are the same !!

OpenStudy (aravindg):

take a look at the graph

OpenStudy (anonymous):

yes I see...so there's something I'm doing wrong in my math here !

OpenStudy (anonymous):

u were right they are the same line!

OpenStudy (aravindg):

take a look at the question you have posted !!

OpenStudy (aravindg):

:P yep makes sure u type the qns right :)